please help!!
You go to the bank and get change for $100 in $1 bills and $10 bills and there are 55 bills total. How many $1 bills and how many $10 bills.

Answer by mananth(16946)

let $1 bill be x
$10 bill will be 55-x
..
1*x+10(55-x)=100
x+550-10x=100
-9x+550=100
-550 -550
-9x=100-550
-9x=-450
/-9
x=-450/-9
x= 50 .... $1 bills
y = 5 $10 bills
...
m.ananth@ hotmail.ca
...
you can also solve by a linear system
x+y =55
x+10y=100
...
multiply first equation by -1
-x-y=-55
add it to second equation
9y = 45
/9
y =5 ... $10 bills
x = $1 bills
...